OpenStudy (anonymous):

What is a unit vector that is perpendicular to both V = 1i+3j+2k and W = 4i+2j+1k?

OpenStudy (anonymous):

think we have to compute the cross product and then normalize right?

OpenStudy (anonymous):

you know how to get the cross product?

OpenStudy (anonymous):

yup, is it just that divided by the length of the resultant vector?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

i get \[-1i-7j+10k\] but check it because it is late

OpenStudy (anonymous):

then divide by \[5\sqrt{6}\]

OpenStudy (anonymous):

i got the same, divided by \[\sqrt{150} which reduces \to 5\sqrt{6}\]

OpenStudy (anonymous):

lol, yeah that looks right

OpenStudy (anonymous):

cool thanks for confirming!

OpenStudy (anonymous):

yw

OpenStudy (dumbcow):

just noticed something i think the i term should be positive --> i -7j +10k divided by 5sqrt6 dot product with <1,3,2>should be 0 1-21+20 = 0